A: You invest $20 at the beginning of each month into stocks that are expected to earn 12% per year. How much will your investment be worth in 20 years?

B:

You are needing to borrow money to buy textbooks. Which of the following options is the best choice?

A)

Bank loan with a 19% APR, compounded annually

B)

Credit Card with a 18% APR, compounded monthly

C)

Credit Card with a 18%, APR, compounded daily

D)

Bank loan with 18.5% APR, compounded quarterly

C: You want to have a perpetuity that will pay $10,000 per year forever. You found an investment offering a guaranteed 1.75% per year forever. How much will you need to invest today in order to receive your first perpetuity payment at the end of the year?
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

B 2A. $20.00 Monthly deposit (beginning) APR Monthly interest rate No. monthly invest in 20 years 12% 1.00% 240 6 7 Value of

Cell reference -

A B 1 2 A. Monthly deposit (beginning) 3 APR Monthly interest rate No. monthly invest in 20 years 20 0.12 =C3/12 =12*20 Value

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.